# Audit Cards: Contextualizing AI Evaluations

### v1

@@ Appendix C Audit cards
C.1 Details and justifications of features
This section provides an overview of audit card features, including a concise summary of what to report and why this contextual information matters. A detailed list of aspects for each feature can be found in the checklist in Section C.2.
@@ C.1 Details and justifications of features
C.2 Checklist for audit card
@@ Appendix D Selection of governance documents
    Major jurisdiction: broadly selected by geopolitical and economic centrality, especially in the AI industry (Stanford HAI staff, 2024; Bianzino et al., 2023).
@@
    Direct influence on model evaluations. For model developers, these tended to be their Responsible Scaling Policies or Frontier Safety Frameworks (see METR, a). For government bodies, we selected documents that tracked with more scale and influence (i.e. more national/federal than local, or more binding than not). The EU Code of Practice is a unique outlier in not being strictly binding but in effect being as influential as it enables presumption of conformity with the Act (see European Union and EU monitor ).
@@
    • Brazil: We analyzed the Federal Senate Bill No. 2338/2023 (approved by Senate in 2024, but pending further steps to become law). (Brazilian Federal Senate, 2023)
    • China: Governance of AI models and systems is detailed across various documents, primarily the: Draft AI Law (with scholars comments, translated May 2024, mur ); Internet Information Service Algorithmic Recommendation Management Provisions (with effect from 2022, Creemers et al. (2022)); Measures for the Management of Generative Artificial Intelligence Services (Draft for Comment) 2023 (Huang et al., 2023)); China 2022 rules for deep synthesis (Gu, 2023)); and China (Draft Measure on Ethical Review).
    • EU: We included the keystone AI Act (EU AI Act, 2024) with its accompanying Code of Practice (“CoP”) (European Commission, 2025): The EU AI Act needs to be read in conjunction with the Code of Practice. The Code of Practice is a unique type of document which, while technically not legally binding, enables presumption of conformity with the EU AI Act. In effect, due to the legal uncertainty of not knowing what alternatives would comply, it is likely to be followed by those wishing to be compliant with the EU AI Act, resulting in effectively a binding effect until “harmonized standards” are published.
    • Singapore: We analyzed the Model AI Governance Framework, which is not technically binding but authoritative guidance. Model AI Governance Framework for Generative AI (Singapore, 2020)
    • South Korea: We analyzed the “Basic Act on the Development of Artificial Intelligence and Establishment of Foundation for Trust”, which comes into force Jan 2026. Governance of AI models and systems appears to be done by a keystone AI Act (with effect from 2026) (National Assembly of the Republic of Korea, 2024).
    • US: Given the lack of a US federal-wide AI regulation, we examined President Biden’s Executive Order (The White House, 2023), which provided the leading signal as to how the US might think about governing model evaluations and risk assessments. It is now repealed, but part of its legacy remains with the highlighted importance of technical assessments and the NIST AI Risk Management Framework.
@@
    • Anthropic: Responsible Scaling Policy (Anthropic, 2024)
    • Google DeepMind: Frontier Safety Framework v2 (Google, 2025)
    • Meta: Frontier AI Framework v1.1 (Meta, 2025)
    • Microsoft: Frontier Governance Framework (Microsoft, 2025)
    • OpenAI: Preparedness Framework Beta (OpenAI, 2023)
    • xAI: Draft Risk Management Framework (xAI, 2025)
Appendix E Features underspecified by current EU Code of Practice stipulations about the Model Report
The EU Code of Practice (“CoP”), if ratified, will be the default way to comply with the AI Act [^12] In theory, model providers can showcase compliance through other means that need to be shown to be equivalent to the Code of Practice. However, given legal uncertainty, this presents a less likely avenue for model developers.. As of its Third Draft version, the Code’s Safety and Security section Commitment II.8 details what Model Reports to be submitted to the AI Office are to include. It covers many aspects, including what our paper calls the core principles and specific features (expertise and background of auditors, scope, methodology, and resources involved in model evaluations). However, it does not specify that the Model Report itself must contain information about channels for external review and feedback for the report, or integrity (conflicts of interest) about the process behind the creation of that report.
These exclusions may be understandable due to varying demands and industry concerns of heavy-handedness (Haeck, 2025). Still, it is worth noting as it allows a clear conversation of whether these issues should be codified, where that code should live, and whether more jurisdictions should adopt this approach (Bradford, 2020). These views will have to be updated with regulatory shifts.
Feature: Review and communication While the AI Office is to have access to key pieces of information like the Model Report, unredacted frameworks, models and evaluation results, channels for non-AI Office peer review are not spelled out. There are multiple candidates for who counts as a relevant “peer” (other model auditors, companies, researchers), and issues, including who would be a trusted coordinator of these interactions, different actors’ contexts and timelines for work, preventing test data leakages and evaluation-gaming. Ironing out the difficulties of this implementation is the subject of future work.
Feature: Integrity (Conflicts of interest) It appears to require appropriate checks for conflicts of interest (Safety and Security section glossary definition of “independent external”) when external auditors are involved, but not the explicit reporting of this in the submitted Model Report.
While we have not included this as part of our audit cards, we note other experts calling for good publishing norms for evaluation reports, which includes: evaluations must be reported per risk and in relation to any risk identification and mitigation efforts by the developer, why certain statistical validity measures were chosen, and how meaningful are the differences between newly tested versus older models or human baselines (see interpretation and quantitative results in Reuel et al. (2024b); Salaudeen et al. ). This non-conclusive list are ways to help others understand why an evaluation finding matters, making it harder to take an evaluation out of context, and easier for a wider community to understand existing coverage of risk assessments, plan for future assessments and how to improve them.
Appendix F Stakeholder interviews
F.1 Interviewee details
@@ F.2 Interview guide
    • Audit cards are a way to standardize reporting of AI evaluation process, with the goal of contextualizing what evals can do and transparency about the process
@@
    • Contribution: Used to color in details about evals process from stakeholders.
@@
    Describe the work you do in relation to evals, try to be specific about your responsibilities, identities of people/orgs you relate to, your role in the evals field
@@
    In your view, what are the biggest challenge preventing evals from being better? More useful for improving safety, or whatever their key goal is?
@@
    Do you consider the limitations and assumptions of your evals process? Where does that thinking get captured (if it does)?
@@
    Can you give us examples of evals where the evals process differed and whether it affected the quality of eval?
@@
      What’s the right balance of transparency in reporting evals, and what risks surround the achieving of that?
@@
    Do you consider the limitations and assumptions of your evals? Where does that thinking get captured (if it does)?
@@
      Do you include different stakeholders in the evals process—who are they, and how are they engaged?
@@
    What are the conditions of you being able to access them? E.g. Recent Cyber CBRN Agent evals report used blue/red anonymized models at UK AISI
@@
    How do you understand evals (as distinct from audits) and what type of evaluation reporting/communication are you aware of?
@@
    What insights are people most often seeking from evals, and how are evals (not) meeting that need?
@@
    Re target audience question: Why do you think [previous answer] is the relevant audience? Other relevant audience segments, how does the messaging change? How can evals most strongly communicate their goal?

### v2

@@ Appendix C Audit cards
Details and justifications of features
This section provides an overview of audit card features, including a concise summary of what to report and why this contextual information matters. A detailed list of aspects for each feature can be found in the checklist in Appendix C.
@@
Checklist for audit card
@@ Appendix D Selection of governance documents
    Major jurisdiction: broadly selected by geopolitical and economic centrality, especially in the AI industry (Stanford HAI staff 2024; Bianzino et al. 2023).
@@
    Direct influence on model evaluations. For model developers, these tended to be their Responsible Scaling Policies or Frontier Safety Frameworks (see METR a). For government bodies, we selected documents that tracked with more scale and influence (i.e. more national/federal than local, or more binding than not). The EU Code of Practice is a unique outlier in not being strictly binding but in effect being as influential as it enables presumption of conformity with the Act (see European Union and EU monitor ).
@@
    • Brazil: We analyzed the Federal Senate Bill No. 2338/2023 (Brazilian Federal Senate 2023). It was approved by Senate in 2024, but still pending approval by the House of Representatives and President to become law.
    • China: Governance of AI models and systems is detailed across various documents. This analysis accounts for: Security Specification for Generative AI Pre-training and Fine-tuning Data (State Administration for Market Regulation and Standardization Administration of China 2025), Draft AI Law, translated with scholars comments (Drafting Expert Group 2025); Internet Information Service Algorithmic Recommendation Management Provisions with effect from 2022 (Creemers, Webster, and Helen Toner 2022); Measures for the Management of Generative Artificial Intelligence Services (Draft for Comment) 2023 (Huang et al. 2023); China 2022 rules for deep synthesis (Gu 2023); and China (Draft Measure on Ethical Review).
    • EU: We included the keystone AI Act (EU AI Act 2024); with its accompanying Code of Practice (“CoP”) (European Commission 2025). The EU AI Act needs to be read in conjunction with the Code of Practice. The Code of Practice is a unique type of document which, while technically not legally binding, enables presumption of conformity with the EU AI Act. In effect, due to the legal uncertainty of not knowing what alternatives would comply, it is likely to be followed by those wishing to be compliant with the EU AI Act, resulting in effectively a binding effect until “harmonized standards” are published.
    • Singapore: We analyzed the Model AI Governance Framework, which is not technically binding but authoritative guidance. Model AI Governance Framework for Generative AI (Singapore 2024).
    • South Korea: We analyzed the “Basic Act on the Development of Artificial Intelligence and Establishment of Foundation for Trust”, which comes into force Jan 2026. Governance of AI models and systems appears to be done by a keystone AI Act with effect from 2026 (National Assembly of the Republic of Korea 2024).
    • US: Given the lack of a US federal-wide AI regulation, we examined the Trump Administration’s AI Action Plan issued in July 2025. This does not have the status of binding law but is the most authoritative signal at the federal level as to how the US government might think about governing model evaluations and risk assessments (The White House 2025).
  • Other norms
    • Bletchley: The Bletchley Declaration by Countries Attending the AI Safety Summit, 1-2 November 2023 (UK Government 2023)
    • Int’l Report: International AI Safety Report (UK AI Safety Institute 2025)
    • Japan AISI: Guide to Evaluation Perspectives (Japan AI Safety Institute 2024)
    • METR: Key Components of an RSP (METR 2024)
    • NIST AI RMF: AI Risk Management Framework (National Institute of Standards and Technology 2024)
    • Paris: Statement on Inclusive and Sustainable Artificial Intelligence for People and the Planet (Élysée Palace 2025)
    • Seoul: Seoul Declaration for Safe, Innovative and Inclusive AI (UK Government 2024)
    • UK AISI: AI Safety Institute Approach to Evaluations (UK AI Safety Institute 2024)
@@
    • Anthropic: Responsible Scaling Policy (Anthropic 2025a)
    • Google DeepMind: Frontier Safety Framework v2 (Google 2025)
    • Meta: Frontier AI Framework v1.1 (Meta 2025)
    • Microsoft: Frontier Governance Framework (Microsoft 2025)
    • OpenAI: Preparedness Framework Beta (OpenAI 2025b)
    • xAI: Draft Risk Management Framework (xAI 2025)
Appendix E Stakeholder interviews
Detailed Methodology
We engaged with stakeholders through interviews in two stages. During literature review and feature selection (see Section 3), we engaged in a dozen off-the-record discussions with academics, evaluation report writers, and policy researchers. After completing the audit card framework, we engaged in the ten semi-structured interviews with stakeholder experts (see Section 6). Our protocol focused on current practices, perceived gaps, and contextual needs.
During both stakeholder engagement stages, the feedback shaped our audit card design. For instance, we distinguished compensation from conflicts of interest (“Integrity”), or condensed features within “Review and feedback”. Interviewees also generally agreed with the high-level value of this audit card proposal, but some expressed doubts about implementation, resulting in the addition of Appendix C.
The interview process was fully based on informed consent. We implemented a comprehensive consent process, obtaining explicit agreement on recording, attribution levels, and data usage. Participants chose their preferred anonymity level, and all quotes were verified before inclusion. Participants were not compensated for the 30-minute voluntary interviews.
Our procedure for reaching out to interviewees was as follows:
As part of the [anonymized for review], we are working on a research project on reporting contextual aspects of the AI evaluations process. The goal is to improve the quality of evaluations through clearer norms on responsible reporting. Here is a more detailed project summary, and below you find the types of information an extensive literature survey has shown to be relevant to evaluations reporting.
As part of this project, we are now looking to get the input from diverse stakeholders in the evaluations process. This could either be in writing or during a short 30-minute call. We would be grateful if you could answer a few questions, which, depending on your exact responsibilities as [role of interviewee], would be some subset of the following:
[Questions as shown in Appendix E.]
Any contribution of yours will be appropriately cited following academic standards. Please let me know if you have any preferences regarding direct quotation or paraphrasing. We are happy to adjust attribution to reflect your preferred level of visibility, whether by name, role at [interviewee organisation], or role more generally. We are happy to share any citations or attributions with you before publication to ensure accuracy and alignment with your preferences.
Finally, we followed up with interviewees to confirm the level of attribution:
We’d be very grateful if you could let us know by Wednesday EOD if you are alright with how this information is being presented. If it’s not acceptable to you, we’d be grateful to understand why and how we can better frame it for better accuracy or anonymity. Any changes to the final draft will be in the direction of less detail/identifiability.
This is how your identity is described: [description of interviewee as per Table 5.]
This is how your interview is being attributed (you are $P_{x}$], highlighted is where you’re cited alone, $P_{y}-P_{z}$ indicate other interviews being referenced together with yours):
[Selection from Section 6 that references the interviewee.]
Interviewee details
See Table 5 for details on the background of interviewees.
@@ Appendix E Stakeholder interviews
    • Audit cards are a way to standardize reporting of AI evaluation process, with the goal of contextualizing what evaluations can do and transparency about the process
@@
    • Contribution: Used to color in details about evaluations process from stakeholders.
@@
    Describe the work you do in relation to evaluations, try to be specific about your responsibilities, identities of people/orgs you relate to, your role in the evaluations field
@@
    In your view, what are the biggest challenge preventing evaluations from being better? More useful for improving safety, or whatever their key goal is?
@@
    Do you consider the limitations and assumptions of your evaluations process? Where does that thinking get captured (if it does)?
@@
    Can you give us examples of evaluations where the evaluations process differed and whether it affected the quality of eval?
@@
      What’s the right balance of transparency in reporting evaluations, and what risks surround the achieving of that?
@@
    Do you consider the limitations and assumptions of your evaluations? Where does that thinking get captured (if it does)?
@@
      Do you include different stakeholders in the evaluations process—who are they, and how are they engaged?
@@
    What are the conditions of you being able to access them? E.g. Recent Cyber CBRN Agent evaluations report used blue/red anonymized models at UK AISI
@@
    How do you understand evaluations (as distinct from audits) and what type of evaluation reporting/communication are you aware of?
@@
    What insights are people most often seeking from evaluations, and how are evaluations (not) meeting that need?
@@
    Re target audience question: Why do you think [previous answer] is the relevant audience? Other relevant audience segments, how does the messaging change? How can evaluations most strongly communicate their goal?